Delta x is given Vo=50 m//s Delta x=100 m Vf= 0m/s Find acceleration and time for plane to drop? (Where would I go about to start this problem?)
Answer #1

I'm assuming these are constant acceleration problems. So to solve these type of question we have to use kinematic equation:

Using 3rd kinematic equation

Vf^2 = V0^2 + 2*a*delta x

Using given values:

a = (Vf^2 - V0^2)/(2*delta x)

a = (0^2 - 50^2)/(2*100)

a = acceleration = -12.5 m/sec^2

Now using 1st kinematic equation:

Vf = V0 + a*t

t = time for plane to drop = ?

t = (Vf - V0)/a

t = (0 - 50)/(-12.5)

t = time taken = 4 sec

Please Upvote.

Comment below if you have any query.